Circular No. 8442

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2004gc IN ARP 327 Independent discovery reports have been received of an apparent su-

pernova in Arp 327 by A. del Olmo, A. Martinez, S. Pedraz, and M. Alises at Calar Alto (2.2-m telescope + CAFOS focal reducer; B, V, R, and I images) on Nov. 18.106 and 19.014 UT and by O. Trondal and M. Schwartz (cf. IAUC 8419; Tenagra II 0.81-m telescope, unfiltered CCD images) on Nov. 18.4 and 19.2 (at mag ∼ 17.4). Schwartz provides the following position for the new object: α = 5^{h}21^{m}49^{s}.95, δ = +6^{o}40′33′′.7, which is 2′′.0 east and 3′′.1 south of the galaxy’s nucleus. Nothing is visible at this location on a Tenagra II image taken on 2002 Oct. 26.4 (limiting mag ∼ 19.5).

SUPERNOVA 1999bw IN NGC 3198 B. Sugerman and M. Meixner, Space Telescope Science Institute; and

J. Fabbri and M. Barlow, University College, London, report the detection of the type-IIn supernova 1999bw in archival Spitzer IRAC images of NGC 3198 obtained by the SINGS Legacy program on 2004 May 1.4 UT. A source was detected in all four IRAC bands at α = 10^{h}19^{m}46^{s}.80, δ = +45^{o}31′35′′.4 (equinox 2000.0; ± 0′′.3 in each coordinate), in close agreement with the optical position (IAUC 7149). The measured flux densities (mJy) at 3.6, 4.5, 5.8, and 8.0 μm are 0.02 ± 0.01, 0.04 ± 0.01, 0.11 ± 0.02, and 0.19 ± 0.04, respectively. This rise is fitted by a 450-K blackbody with integrated flux 3.1 × 10^{−13} erg cm^{−2} s^{−1}, which, for a distance to NGC 3198 of 14.5 Mpc (Kelson et al. 1999, Ap.J. 514, 614), corresponds to a luminosity of 6.2 × 10^{38} ergs s^{−1} and a blackbody radius of 1.6 × 10^{16} cm. This size is consistent with ejecta expanding at 1000 km/s in the five years since core collapse, suggesting the reported emission may be from dust that condensed within the ejecta. A convolved companion source is also detected, separated by 3′′.9 at p.a. 209^{o}, with 3.6-, 4.5-, 5.8-, and 8.0-μm fluxes of 0.08 ± 0.01, 0.07 ± 0.01, 0.33 ± 0.04, and 0.88 ± 0.08 mJy, respectively.

SUPERNOVA 2004et IN NGC 6946 Visual magnitude estimates: Sept. 29.774 UT, 12.8 (K. Hornoch, Lele-

kovice, Czech Republic); Oct. 3.86, 12.9 (J. Carvajal, Madrid, Spain); 11.908, 12.8 (Hornoch); 15.483, 12.4 (S. Yoshida, Gunma, Japan); Nov. 9.490, 12.7 (Yoshida).

2004 November 20  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green